OpenStudy (ccoffey415):

(x^2-6x+5)-(2x^2+x+1)

OpenStudy (photon336):

first distribute the negative sign: then combine like terms (x^2-6x+5)-(2x^2-x-1) (x^2)-(2x^2) = -x^2 -6x-x = -7x -x^2-7x-4